Presentation is loading. Please wait.

Presentation is loading. Please wait.

K. Long, 18 April, 2015 Beam-line status — red sky at night?

Similar presentations

Presentation on theme: "K. Long, 18 April, 2015 Beam-line status — red sky at night?"— Presentation transcript:

1 K. Long, 18 April, 2015 Beam-line status — red sky at night?

2 Upstream beam line: – Steel shielding removed – Q1—3, D2 re-instated – Soak test performed

3 Decay solenoid:


5    ‘In hand’: CMacW assisting sol d team to finalise QD While some documentation exists, thorough job to generate consistent set yet to be completed M.Courthold, M.Hills Presented to and agreed by MICE Technical Board

6 Downstream beam line & Step I instrumentation: D2, Q4—Q9: – Water and power re- established Pressure reducing valves fitted – Need to be calibrated Magnets have been ‘dry’ for weeks Soak test OK Beam stop – Works OK Step I instrumentation: (installation) – Inside DSA complete: GVA1, BM1, TOF0, Ckov – Outside DSA: BM2, TOF1 in position KL to be positioned

7 D1 tested for 2h10min at 245 A and for 30min at 400 A Q1-2-3 tested at 195 A for 3h Q4-5-6-7-8-9 tested at 395 A for 6h D2 tested at 150A for 5h40m and at 120A for 30m DKsol tested at 60A for >3h temp. taken at different intervals showed stability Some MPS+MAGNET waterflow failures resulting in trips: however no evidence of water flow problems was found CONCLUSIONS –Confident the line is fine –shame no time to test D1 at 400A for longer period –DKsol tested separately at 870A –Important to sort out the fake waterflow issues Q1-2-3 D1 DKsol D2 Q4-5-6-7-8-9 Soak Test - August 26 th 2009 (MA, A.Dobbs, M.Ellis, K.Long)

8 BREAKDOWN of Q1-2-3 D1 DKsol D2 Q4-5-6-7-8-9 Soak Test - August 26 th 2009 In the morning (10:00 to 13:20) only DS BL was up. US BL was not powerable due to hardware intervention (soldering, M. Hughes) DS BL morning settings: D2=195A Q4-5-6-7-8-9=395 A At 13:30 D2 was set off to reduce overall power and allow US beamline to be tested Q1-2-3 = 195 A D1=245 A  value chosen to not overcame an alleged software threshold DKsol=60A D2= 0A Q4-5-6-7-8-9=395A We experienced some MPS and MAGNET waterflow failures which triggered interlocks. In all cases this did NOT seem to correspond to real failures rather some mis-reading from the unit. This should be understood or we can risk abrupt stops when running in Sept/Oct D1 tripped @ 13:40  set back to 245A 14:20: D2 set to 150 A D2,Q4,Q6 tripped @ 15:20  set back to previous values 15:40 Q1 Q2 Q3 D1 DK D2 Q4 Q5 Q6 Q7 Q8 Q9 195 195 195 245 60 150 395 395 395 395 395 395 16:00 set D1 to higher value (no software threshold preventing it !) 195 195 195 400 60 120 200 280 200 200 280 350  the final configuration (30 min test) Temperature readings taken at different times after several hours showed a good stability CONCLUSION: we believe magnets are well behaving.We tested the line with all magnets at values close to their maximum. We could not power them all MAX simultaneously so We still need to understand the MPS/MGNET waterflow failures generating interlocks which do not seem to correspond to REAL events

9 Target Just pictures, details from C.Booth

10 Target build: Shaft received from TechVac (DLC coating company) Wednesday 19Aug09 Assembly of ‘Target 1’ (T1) complete Friday 21Aug09 – Installation in R78 completed Saturday 22Aug09 – 50k actuations completed Sunday 23Aug09 – Metrology of target tip position within dummy vacuum vessel Monday 24Aug09 – Installation in ISIS complete Tuesday 25Aug09 Gate valve opened ~16:30 – Target space is under the ISIS machine vacuum! Assembly of T2 complete Sunday 23Aug09 – Installation in R78 complete Tuesday 25Aug09

11 Modifications in synchrotron vault: Modified gate valve fitted – Including thermocouple ‘Collection pot’ and mirror system fitted Connections reversed for quick installation and removal

12 T1 build:

13 T1 installed on ISIS:

14 T2 build and installation in R78: Great team effort! – Great support from ISIS personnel Outstanding issues: – See C.Booth talk, but headlines: Details of controls and EPICS interfaces in MLCR installation being sorted out Debugging non-standard actuation behaviour in R78 set up Still to perform 400k actuation run in R78

15 Conclusions: Magnets that make up MICE Muon Mean ready for operation Target installed and being debugged Offline target installed and in commissioning phase: – Debugging non-standard actuations ‘Rehearsal’ of running full beam line scheduled for Thursday 03Sep09 For next MICE running phase: – MMB expert on call 24/7: Expert to be present to assist MOM and shift crew in setting up beam line for data taking and to shutdown beam line at the end of data taking: – Essential: e.g. one target interlock not yet established Rota to be established: – MMB expert phone obtained – MOM/shift should call MMB expert first – Goal of all this to be more efficient in set-up, operation, and shutdown than last year And finally: – MICE shutdown briefing: 12:00, 04Sep09, MICE Hall

Download ppt "K. Long, 18 April, 2015 Beam-line status — red sky at night?"

Similar presentations

Ads by Google